There are six mutually exclusive investments alternatives. And investor can select one of them or non .the following analysis table shows incremental internal rate of return for each two alternatives
(2.5 points) There are six mutually exclusive investment alternatives. An investor can select one of them or none. The following analysis table shows incremental internal rate of retum for each two altematives. (e-t. incremental internal rate of retum of Δ(B-A)-8%], which alternative should he select if MAR-12% Do Nothing A Data Initial Investment Do Nothing $10,000 | $11,000-12,000 | $14,000 | $15,000| $17,000 14% í 11% i 12% 14% 13% 13% 15% 14% 13% 8% 8% 13% 11% 11% 7% 15% 7% 10% 14% a) Altemative A d) Altemative D g) Do nothing altenative b) Altermative B e) Altemative B c) Alternative C 1) AltemativeF ) Not enough information
Answer #1

Out of these 6 alternatives one of the IRR is greater than MARR therefore reject the Do nothing alternative.

Now, determine the lowest invest alternative here alternative A has lowest initial cost. Current selection is A

Now, compare A with next best alternative that is B

Incremental RoR of B - A = 8%

Since it is less than MARR (12%) therefore select the least expensive alternative. Here alternative A will be selected.

Now compare A with 2nd best alternative that is C

large riangle (C -A) = 14%

Since it is greater than MARR therefore reject A and select C.

Now compare C with 3rd best alternative that is D

large riangle (D - C) = 13%

Since greater than MARR therefore reject C and select D.

Now compare D with E

NE-D) = 7%

Less than MARR therefore reject E and select D.

Now compare D with F

large riangle (F - D) = 10%

Less than MARR therefore reject F and select D.

Finally D must be selected.

D. Alternative D

Please contact if having any query thank you.